# Improving Dialog Systems with Natural Language Feedback and Preferences

### v1

@@ 5 Case Study 1: Query Generation
5.1 Derive Criteria from Feedback
We collect all natural language feedback from the FITS training split to understand human preferences and derive criteria. We first use SimCSE sentence encoder555 We use the sup-simcse-roberta-large model. (Gao et al., 2021) to encode each feedback. Then we use $k$-means clustering to group all feedback related to query generation into five clusters and inspect the feedback in different clusters. Overall, we summarize the following four groups (see Table 1) of feedback types through merging clusters. By analyzing Table 1, we derive that a successful search query should (i) rephrase the user’s question while keeping important keywords, (ii) be relevant and specific, (iii) use some common words for better search coverage when possible, (iv) be concise. The exact criteria text which will be used to construct the prompt $c$ for the text refiner $P_{\phi}(r^{\prime}|r,q,c)$ is given in Table 2.
5.1.1 Criteria-guided Metric Design
After deriving criteria for improved query generation from the supplied feedback, we then design the following metrics to measure aspects of the quality of a search query that users indicated they care about in the feedback.666 When measuring the performance of a set of queries, for a metric defined as a fraction with a constant numerator, we take the average of the denominators of all queries on that metric and take its reciprocal to multiply the numerator. These metrics are designed with the guideline that a good search query will score highly in all metrics in a balanced way.
Non-copy rate
measures how much the search query rephrases the user’s utterance by examining $n$-gram matching. In particular, we define “Non-copy Rate” in Equation 1, which is based on BLEU-4 (Papineni et al., 2002). In the equation, $s$ is the search query and $u$ is the user question.
[…]
Specificity
measures whether the search query is specific enough to cover all the information needed to be able to search for the most relevant documents. We use GPT-3.5 to measure specificity (Fu et al., 2023) where we concatenate the dialog context and search query together and ask GPT-3.5 to judge whether the search query is specific using the chain-of-thought technique (Wei et al., 2022). A meta-evaluation shows 80% agreement between human annotators and GPT-3.5 for evaluating query specificity. Details can be found in the Appendix.
Readability
measures how readable is the search query by examining the word frequency rank (WFR)777 We use the list of most frequent 333,333 words in English provided by Kaggle: https://www.kaggle.com/rtatman/english-word-frequency of words in the search query. Intuitively we want the search query to contain some common words to enhance readability. We define “Readability” in Equation 2. In the equation, $w$ represents a word in $s$, $C$ is a large constant for rescaling so that the readability numbers will not be too small.
$\text{Readability}=\frac{C}{\text{AVG}_{w\in s}(\text{WFR}(w))}$ (2)
Conciseness
measures how concise the search query is by counting the number of words in the search query. We use the reciprocal of the length of the query to denote “Conciseness”. We also multiply by a constant 100 for rescaling to make a similar scale to other metrics.
Coverage
measures how specific vs. general a search query is by counting the number of Google search result pages. Since the number of pages is highly variable, we define this metric as a relative one. Given the same dialog context and the refined queries obtained using Table 2, only the query that gets the most search results has “Coverage” of 1, other queries will have “Coverage” of 0.
Satisfaction
measures whether the search query will satisfy the user. It is an overall metric, and we use our trained satisfaction classifier to measure the percentage of satisfied refinements.
5.2 Construct refinement training data
Prompt for query refinement with GPT-3.5
[Criteria]
Below is the dialog context.
[Dialog Context]
Below is the bot’s unsatisfactory query.
[Original Query]
You should modify the original search query into the following:
Valid Test Test Unseen
F1 PPL F1 PPL F1 PPL
BB2(QG) 9.74 16.09 14.28 9.61 16.09 10.15
[…]
Table 4: Performance of different query generators on FITS validation, test, and test unseen datasets, evaluated using standard F1 and PPL. The baseline BlenderBot 2 method is outperformed by our system-level trained query generator using both satisfied and refinement data (last row). Using only satisfied data is not as good (middle row).
Valid Test Test Unseen
[…]
Table 5: Performance of different query generators on FITS validation, test, and test unseen datasets, evaluated using our designed metrics. NCR: non-copy rate, Spec.: specificity, Read.: readability, Con.: conciseness, Cov.: coverage. Sat.: satisfaction. The original BlenderBot 2 generator (top row) generates queries that have low satisfaction, despite scoring well in some other metrics. Our system-level trained query generator using both satisfied and refinement data (last row) provides much better satisfaction. Using only satisfied data is not as good (middle row).
Group Feedback type Number Percentage
1
Clarify his/her demand again.
3702 26.54%
2
Complain that the bot (i) does not answer the question (ii) or gives irrelevant information (iii) or asks the user to find out the answer on his or her own.
2260 16.20%
3
Point out specific search results that can answer the question.
2255 16.17%
4
Suggest that the bot should use the search results.
2130 15.27%
5
Point out that the bot’s answer is (i) factually incorrect, (ii) or not grounded on the search results.
1572 11.27%
6
1309 9.39 %
7
582 4.17%
8
Complain about repetition/rudeness in bot responses.
137 0.99%
Table 6: Case study 2 (response generation): Eight groups of system-level feedback derived from automatic clustering.
We first randomly sample 1,000 satisfied queries from the FITS training set together with their contexts to add to our supervised training data $\mathcal{D}$. Then, according to Figure 1-(B), for each unsatisfied Internet search query $r$, we go through the following three steps (1) refinement generation: use GPT-3.5 in combination with the criteria $c$ derived from §5.1 to get a refinement $r^{\prime}$. (2) Quality check: use a quality checker to check whether $r^{\prime}$ is satisfactory. (3) Collect data: add $(q,r^{\prime})$ to the training data $\mathcal{D}$ if $r^{\prime}$ is satisfactory, otherwise, discard it. We will describe the process of (1) in detail in the following section.
5.2.1 Refinement Generation
We instruct GPT-3.5 to generate refinements for 1,000 randomly sampled unsatisfied Internet search queries through criteria-based prompts, as shown in Table 3. To demonstrate the effectiveness of Figure 1-(A), we conduct ablation studies of using different criteria to prompt GPT-3.5 for query refinement. Considering our computational budget, for metrics calculated using GPT-3.5, we randomly sampled 500 dialog contexts and compare the queries resulting from different criteria.
The results are given in Table 2. It is clear that adding criteria in the prompt will shift the generation of GPT-3.5, and the performance differences are interpretable using the newly designed metrics. We make the following observations: (i) The rephrase criterion increases the non-copy rate. (ii) The relevance criterion increases the relevance metric. (iii) The readability criterion increases the readability and coverage metrics. (iv) When we add all the criteria, the refinements can achieve reasonably good performance in all our designed perspectives as well as in overall satisfaction. Therefore, when collecting training data, we use the four criteria augmented prompt for query refinement.
5.3 Fine-tuning the Model
In our experiments, we used the 400M parameter BlenderBot 2 query generator based on BART (Lewis et al., 2020). We consider two fine-tuning settings: (1) one with only the satisfied data; and (2) with both satisfied and refinement data. During training, we use the Adam optimizer (Kingma and Ba, 2015) with a batch size of 8 and learning rate of $7\times 10^{-6}$ for three epochs, and the best model checkpoint is chosen based on validation loss.
5.4 Evaluation using designed metrics
We compare the performance of the following query generators using both standard metrics and our designed metrics.
BB2(QG)
The original BB2 query generator.
[…]
Our system-level trained query generator using satisfied data only.
[…]
Our system-level trained query generator using both satisfied and refinement data.
5.4.1 Results on Standard Metrics
Results using standard metrics (following Shi et al. (2022)) are shown in Table 4. Comparing the performance of the original BB2 query generator and the one trained with satisfied data, it is clear that training the query generator with domain-specific data helps improve performance on F1 word overlap and perplexity (PPL) metrics by a large margin. In addition, adding refinement data for fine-tuning further enhances the performance of both metrics.
5.4.2 Results on Our Designed Metrics
We also report results on our designed metrics for different query generators, as shown in Table 5. It is clear that by training on satisfied data, the generated queries are more specific and accurate, while further adding refinement data for training will lead to even better specificity. The original BB2 query generator generates queries that are too short to be used to retrieve the most relevant documents regarding the user’s question. In other words, although it generates queries that perform well in terms of readability or coverage, it is still an inadequate query generator, as evidenced by the poor satisfaction of the queries it generates. Later, when we refer to “our trained query generator”, we mean the one trained using both satisfied data and refinement data.
6 Case Study 2: Response Generation
@@
6.1 Derive criteria from feedback
Using a similar method to §5.1, we group all feedback related to response generation into ten clusters. Then, we summarize the following eight groups (see Table 6) of feedback types through splitting or merging different clusters. By analyzing Table 6, we derive that an improved response as indicated by users should (i) ground its answer on relevant search results, (ii) be concise and targeted, (iii) be confident in its answer. The exact criteria text which will be used to construct the prompts $c$ for the text refiner $P_{\phi}(r^{\prime}|r,q,c)$ is given in Table 7.
6.1.1 Criteria-guided Metric Design
After deriving criteria for response generation from feedback, we design the following metrics to measure the quality of a response as indicated by users.888 When using a following metric to measure the performance of a set of responses, we simply take the average of all responses’ scores on that metric.
Groundedness
measures how much the response utilizes the search results by examining $n$-gram matching. We define “Groundedness” in Equation 3, which is based on ROUGE-2 (Lin, 2004). In the equation, $r$ is the final response, and $d$ is a document from a set of relevant search documents $\mathcal{S}$.
$\text{Groundedness}=\max_{d\in\mathcal{S}}\text{ROUGE-2}(r,d)$ (3)
Factuality
measures whether the information in the response is supported by search documents. We use GPT-3.5 with chain-of-thought reasoning to measure factuality (Luo et al., 2023). We concatenate the search documents and response, and ask GPT-3.5 to judge if all information in the response is supported by the search documents. A meta-evaluation shows 88% agreement between human annotators and GPT-3.5 for evaluating response factuality. Details of the prompts and meta-evaluation can be found in the Appendix.
Helpfulness
measures whether the response directly answers the user’s question. Similar to the measurement of factuality, we use GPT-3.5 to measure helpfulness. The details are in the Appendix.
Relevance
measures whether the response provides pertinent information and stays on track without deviating from the subject. We again use GPT-3.5, with further details in the Appendix.
Confidence
measures whether the response is in a certain and confident tone. We use simple heuristics to gauge response confidence, counting the occurrences of “I’m not sure” and “I don’t know.” If either phrase appears, we consider the response as unconfident; otherwise, we assume the response to be confident.
Satisfaction
measures whether the response satisfies the user, similar to “satisfaction” in §5.1.1 .
6.2 Construct refinement training data
Prompt for response refinement with GPT-3.5
[Criteria]
Below is the dialog context.
[Dialog Context]
Below is the bot’s unsatisfactory response.
[Original Response]
Below are some useful search results that you could use.
[Search Documents]
You should modify the original response into the following:
Valid Test Test Unseen
F1 PPL F1 PPL F1 PPL
BB2(QG+RG) 25.78 9.40 28.30 7.41 22.99 7.75
SLT(QG)+BB2(RG) 26.69 8.24 28.66 6.66 24.88 7.03
[…]
Table 9: Case study 2 (response generation): performance of different dialog systems on FITS validation, test, and test unseen datasets, evaluated using standard F1 and PPL. The original BlenderBot 2 system (top row) is outperformed by our system-level trained response generator using satisfied data (third row). However, using refinement data in this case (last row) did not help, see §6.4.1 for further anaslysis.
Valid Test Test Unseen
[…]
Table 10: Case study 2 (response generation): performance of different dialog systems on FITS validation, test, and test unseen datasets, evaluated using our designed metrics. GRD: groundedness, Fact.: factuality, Help.: helpfulness, Rel.: relevance, Conf.: confidence. Sat.: satisfaction. The original BlenderBot 2 system (top row) is outperformed by our system-level trained response generator using satisfied and refinement data (last row). Using only satisfied data (third row) or only our query generator (second row) is not as good.
Similar to §5.2, we first randomly sample 1,000 satisfied responses together with their contexts to add to our training data $\mathcal{D}$. Then, we go through the following three steps: (1) refinement generation, (2) quality check and (3) collection of filtered data. We will describe (1) in detail in the following section.
6.2.1 Refinement Generation
We instruct GPT-3.5 to generate refinements for 1,000 randomly sampled unsatisfied responses through criteria-based prompts, as shown in Table 8. As in §5.2.1, we conduct ablation studies to demonstrate the effectiveness of derived criteria from feedback. The results are shown in Table 7. We make the following observations: (i) Adding the groundedness criterion will improve the groundedness metric. (ii) Adding the relevance criterion will increase helpfulness and relevance. (iii) GPT-3.5 generated refinements are always confident and rarely include sentences like “I’m not sure” or “I don’t know”. (iv) In terms of satisfaction, the best overall performance is achieved by the prompt with all the criteria added. Therefore, when collecting training data, we use the three criteria-augmented prompt for response refinement.
6.3 Fine-tuning the Model
In our experiments, we used the 400M parameter BlenderBot 2 main model as the baseline response generator. As before, we consider two fine-tuning settings: (1) using only satisfied data; and (2) with both satisfied and refinement data, following §5.3.
6.4 Evaluation using designed metrics
We compare the performance of the following systems:
BB2(QG+RG)
Original BB2 response model equipped with the original BB2 query generator. This is the baseline system we attempt to improve upon.
SLT(QG)+BB2(RG)
The original BB2 response generator equipped with our best system-level trained query generator from §5.
[…]
Our system-level trained response generator using satisfied data only; equipped with our best query generator from §5.
[…]
Our system-level trained response generator using both satisfied and refinement data; equipped with our best system-level trained query generator from §5.
6.4.1 Results on Standard Metrics

### v2

@@ 5 Case Study 1: Query Generation
Type Criteria (Abbreviated) NCR Spec. Read. Con. Cov. Sat.
(1): Baseline None 4.06 79.40 19.46 14.87 29.80 61.50
(2): (1)+Rephrase Rephrase the user’s question and keep keywords. 4.98 83.20 19.54 15.04 26.50 62.10
(3): (2)+Specificity Above + Be accurate and specific for user needs. 5.00 84.20 18.77 14.50 28.80 63.30
(4): (3)+Readability Above + Use simple and common words for better results. 5.08 80.80 19.53 15.97 29.40 62.40
(5): (4)+Conciseness Above + Be concise; focus on user’s first question. 4.81 80.00 19.70 16.63 35.30 62.70
Table 2: Case study 1 (query generation): refinement quality via designed metrics when using different criteria to prompt GPT-3.5 for query refinement. NCR: non-copy rate, Spec.: specificity, Read.: readability, Con.: conciseness, Cov.: coverage. Sat.: satisfaction. The full criteria texts can be found in the Appendix.
5.1 Derive Criteria from Feedback
We collect all NL feedback from the FITS training split to understand human preferences and derive criteria. We first use SimCSE encoder444 We use the sup-simcse-roberta-large model. (Gao et al., 2021) to encode each feedback. Then, we use $k$-means clustering to group feedback related to query generation into five clusters. From inspecting these (see Appendix for detailed manual efforts), we summarize them into four groups (see Table 1) and derive that a successful search query should (i) rephrase the user’s question while keeping important keywords, (ii) be relevant and specific, (iii) use common words for better search coverage, (iv) be concise. The criteria text for crafting the prompt $c$ for the text refiner $P_{\phi}(r^{\prime}|r,q,c)$ is in Table 2.
5.1.1 Criteria-guided Metric Design
Using feedback-derived criteria, we design metrics to mirror users’ preferences.555 When evaluating a set of queries, for a metric defined as a fraction with a constant numerator, we take the average of the denominators of all queries on that metric and take its reciprocal to multiply the numerator. Ideally, an effective query should score high across all these metrics.
  • Non-copy rate measures how much a search query rephrases the user’s utterance by examining $n$-gram matching. We define it in Equation 1 based on BLEU-4 (Papineni et al., 2002) where $s$ is the search query and $u$ is the user question.
[…]
  • Specificity measures whether the search query sufficiently captures the necessary information to retrieve relevant documents. We use GPT-3.5 as the evaluator (Fu et al., 2023). Details are in Appendix.
  • Readability measures a search query’s clarity based on the word frequency rank (WFR)666 We use the Kaggle dataset for WFR: https://www.kaggle.com/rtatman/english-word-frequency of its terms, as defined in Equation 2, where $w$ is a word in $s$ and $C$ is a scaling constant. Ideally, a query should use common words to improve readability.
    $\text{Readability}=\frac{C}{\text{AVG}_{w\in s}(\text{WFR}(w))}$ (2)
  • Conciseness measures the query’s brevity by its word count, with its value being the query length’s reciprocal, scaled by a constant 100.
  • Coverage measures how specific vs. general a search query is by counting the number of Google search result pages. Considering the wide variation in page count, we employ a relative metric. For refined queries obtained using Table 2 with the same dialog context, the query with the most results gets a “Coverage” score of 1, and others receive 0.
  • Satisfaction measures whether the search query will satisfy the user. It is an overall metric, and we use our trained satisfaction classifier to determine the percentage of satisfied refinements.
5.2 Construct refinement training data
We sample 1,000 satisfied queries from the FITS training set along with their contexts to add to our supervised training data $\mathcal{D}$. Then, based on Figure 1-(B), for each unsatisfied query $r$, we (1) use GPT-3.5 and criteria $c$ derived from §5.1 to get a refinement $r^{\prime}$. (2) Use a quality checker to check $r^{\prime}$’s satisfaction. (3) Add $(q,r^{\prime})$ to $\mathcal{D}$ if $r^{\prime}$ is satisfactory. We elaborate on step (1) in the next section.
5.2.1 Refinement Generation
Valid Test Test Unseen
[…]
Table 3: Evaluate query generators on FITS using designed metrics. See Table 2 caption for abbreviation meanings.
Valid Test Test Unseen
F1 PPL F1 PPL F1 PPL
BB2(QG) 9.74 16.09 14.28 9.61 16.09 10.15
[…]
Table 4: Evaluate query generators on FITS using F1 and perplexity (PPL).
We use GPT-3.5 with criteria-based prompts to refine 1,000 randomly sampled unsatisfied queries (details in Appendix). To demonstrate the effectiveness of Figure 1-(A), we conduct ablation studies with different criteria for query refinement. Given our computational budget, for metrics relying on GPT-3.5, we sample 500 dialog contexts and compare the queries resulting from different criteria.
The results are in Table 2. Adding criteria in the prompt will shift GPT-3.5’s generation, and the performance differences are interpretable using our designed metrics. Specifically, (i) The rephrase criterion increases the non-copy rate. (ii) The relevance criterion increases the relevance metric. (iii) The readability criterion increases the readability and coverage metrics. (iv) Using all the criteria, the refinements achieve reasonably good performance in all our designed perspectives and overall satisfaction. Thus, when collecting training data, we use the four criteria augmented prompt for refinement.
5.3 Fine-tuning the Model
We start from the 400M BB2 query generator and consider two fine-tuning settings: (1) using the satisfied data; and (2) using satisfied and refinement data. During training, we use the Adam optimizer (Kingma and Ba, 2015) with a batch size of 8 and learning rate of $7\times 10^{-6}$ for three epochs. The best checkpoint is chosen based on validation loss.
Group Feedback type Num. %
1
Clarify his/her demand again.
3702 26.54%
2
Complain that the bot (1) does not answer the question or (2) gives irrelevant information or (3) asks the user to find out the answer on his or her own.
2260 16.20%
3
Point out specific search results that can answer the question.
2255 16.17%
4
Suggest that the bot should use the search results.
2130 15.27%
5
States that the answer is (1) factually incorrect, or (2) not grounded on the search results.
1572 11.27%
6
1309 9.39 %
7
582 4.17%
8
Complain about repetition/rudeness in bot responses.
137 0.99%
Table 5: Case study 2 (response generation): 8 groups of system-level feedback derived from automatic clustering.
5.4 Evaluation using designed metrics
We evaluate the following query generators.
  • BB2(QG) The original BB2 query generator.
[…]
  • Results on Standard Metrics Table 4 presents the results using standard metrics, as per Shi et al. (2022). Compared to the original BB2 query generator, training with domain-specific data (2nd row) significantly improves F1 word overlap and perplexity metrics. Adding refinement data (3rd row) further enhances these metrics.
  • Results on Our Designed Metrics We also report results on our designed metrics for different query generators in Table 3. It is clear that training on satisfied data produces more specific and accurate queries, with further improvements when incorporating refinement data. The original BB2 query generator often generates overly concise queries, hindering the retrieval of the most relevant documents. In other words, although it generates queries that perform well in terms of readability or coverage, it is still an inadequate query generator, as evidenced by the poor satisfaction of the queries it generates. Later, when we refer to “our trained query generator”, we mean the one trained using both satisfied data and refinement data.
6 Case Study 2: Response Generation
Type Criteria (Abbreviated) GRD Fact. Help. Rel. Conf. Sat.
(1): Baseline Use a conversational tone; no more than 20 words. 34.68 86.60 81.40 89.40 99.60 74.10
(2): (1)+Groundedness Above + Use search results to give answers. 36.81 86.60 85.00 89.00 99.90 75.80
(3): (2)+Relevance Above + Be concise and targeted, no irrelevant information. 36.77 88.80 85.60 89.40 99.90 74.90
(4): (3)+Confidence Above + Don’t start with “I’m not sure” or “I don’t know”. 39.02 87.20 86.60 90.60 99.90 77.00
Table 6: Case study 2 (response generation): refinement quality via designed metrics when using different criteria to prompt GPT-3.5 for response refinement. GRD: groundedness, Fact.: factuality, Help.: helpfulness, Rel.: relevance, Conf.: confidence. Sat.: satisfaction. The full criteria texts can be found in the Appendix.
6.1 Derive criteria from feedback
Following the approach in §5.1, we group all feedback related to response generation into ten clusters. Then, we summarize the following eight groups (see Table 5) of feedback types by merging some clusters. From Table 5, we derive that an improved response as indicated by users should (i) ground its answer on relevant search results, (ii) be concise and targeted, (iii) be confident in its answer. The exact criteria text for crafting the prompts $c$ for the text refiner $P_{\phi}(r^{\prime}|r,q,c)$ is given in Table 6.
6.1.1 Criteria-guided Metric Design
After deriving criteria for response generation from feedback, we design the following metrics to measure the quality of a response as indicated by users.777 When evaluating a set of responses using one of the following metrics, we take the average of all responses’ scores on that metric.
  • Groundedness measures how much the response utilizes the search results by examining $n$-gram matching. We define it in Equation 3 based on ROUGE-2 (Lin, 2004). Here, $r$ is the response, $d$ is a document from the relevant search set $\mathcal{S}$.
    $\text{Groundedness}=\max_{d\in\mathcal{S}}\text{ROUGE-2}(r,d)$ (3)
  • Factuality checks whether the information in the response is backed by search documents. We use GPT-3.5 with chain-of-thought to measure factuality (Luo et al., 2023). See Appendix for details.
  • Helpfulness measures whether the response directly answers the user’s question. We use GPT-3.5 to measure helpfulness. See Appendix for details.
  • Relevance measures whether the response remains on topic and offers pertinent information. We again use GPT-3.5, with further details in the Appendix.
  • Confidence measures whether the response is in a certain and confident tone. We use simple heuristics to gauge confidence, counting the occurrences of “I’m not sure” and “I don’t know.” If either phrase appears, we consider the response unconfident; otherwise, it’s considered confident.
  • Satisfaction measures whether the response satisfies the user, similar to “satisfaction” in §5.1.1.
6.2 Construct refinement training data
Valid Test Test Unseen
[…]
Table 7: Evaluate dialog systems on FITS using designed metrics. See Table 6 caption for abbreviation meanings.
Valid Test Test Unseen
F1 PPL F1 PPL F1 PPL
BB2(QG+RG) 25.78 9.40 28.30 7.41 22.99 7.75
SLT(QG)+BB2(RG) 26.69 8.24 28.66 6.66 24.88 7.03
[…]
Table 8: Evaluate dialog systems on FITS via F1 & PPL.
As in §5.2, we first randomly sample 1,000 satisfied responses together with their contexts to add to our training data $\mathcal{D}$. Then, we go through the following three steps: (1) refinement generation, (2) quality check and (3) collection of filtered data. We will describe (1) in detail in the following section.
6.2.1 Refinement Generation
We use GPT-3.5 with criteria-based prompts to refine 1,000 sampled unsatisfied responses (details in Appendix). As in §5.2.1, we conduct ablation studies to demonstrate the effectiveness of derived criteria. The results in Table 6 highlight: (i) Adding the groundedness criterion improves the groundedness metric. (ii) Adding the relevance criterion increases helpfulness and relevance. (iii) GPT-3.5 refinements are confident and rarely include phrases like “I’m not sure” or “I don’t know”. (iv) In terms of satisfaction, the best performance is achieved by the prompt with all criteria added. Therefore, when collecting training data, we use the three criteria-augmented prompt for response refinement.
6.3 Fine-tuning the Model
We use the 400M BB2 main model as the baseline response generator and consider two fine-tuning settings: (1) using only satisfied data; and (2) using both satisfied and refinement data, following §5.3.
6.4 Evaluation using designed metrics
We evaluate the following systems:
  • BB2(QG+RG) Original BB2 response generator paired with the original BB2 query generator.
  • SLT(QG)+BB2(RG) Original BB2 response model paired with our trained query generator.
[…]
Valid Test Test Unseen
[…]
Table 9: Case study for combining system-level and instance-level feedback: performance of different dialog systems on FITS datasets, evaluated using our designed metrics. See Table 6 for the meaning of the abbreviations.
  • Results on Standard Metrics Standard metrics are shown in Table 8. Key takeaways include: (i) When using the BB2 response generator, our trained query generator improves the final response quality compared to the BB2 query generator. (ii) Training the response generator on satisfied data leads to further improvements when using our best query generator. (iii) However, training with additional refinement data does not surpass using satisfied data alone. The reason behind (iii) relates to FITS’s gold response collection. Often, the gold response is a user-guided, BB2-generated reply. This biases reference-based metrics towards original BB2 outputs. Moreover, low-quality references may underestimate model performance when using reference-based metrics (Zhang et al., 2023a) and we confirmed this with a human evaluation of response quality (see Appendix for details).
  • Results on Our Designed Metrics Table 7 shows the results when using our designed metrics. Notably, (i) when using BB2 response generator, our trained query generator improves the final response quality from all perspectives compared to the BB2 query generator. (ii) When equipped with our trained query generator, training the response generator on satisfied data leads to consistent improvements in helpfulness compared to the BB2 response generator, indicating the importance of domain-adapted training. (iii) Training the response generator on both satisfied and refinement data improves the final response quality from all perspectives compared to training on satisfied data only, highlighting refinement data’s utility in rectifying model errors. (iv) In terms of satisfaction, the best-performing system employs our query and response generators, both trained on satisfied and refinement data. Additionally, we gathered the first 200 unsatisfied responses into a sparse refinement training set, refined via instance-level feedback. A model trained on this set alongside satisfied data, fell short compared to our system-level trained response generator, as measured by our designed metrics (see Appendix for details).
7 Combining System-level Feedback and Instance-level Feedback
Previous studies (Scheurer et al., 2022b; Shi et al., 2022; Chen et al., 2023a) have shown the effectiveness of instance-level feedback in the refinement process. To take a step further, we explore the synergy of system-level and instance-level feedback on dialogue systems. Using response generation as a case study, we collect both human and GPT-3.5 feedback (prompt in Appendix) for the 1,000 unsatisfied responses from §6.2.1. We then design a refinement prompt integrating both system-level and instance-level feedback (see Appendix). We introduce three systems for comparison.
[…]
7.1 Results of Adding Instance-level Feedback
Results using our designed metrics are in Table 9. We observe that adding human-written feedback to the response refinement part brings improvements in the five criteria-based metrics most of the time, and increases the overall satisfaction consistently. However, adding GPT-3.5 feedback results in degraded performance in groundedness, factuality and confidence. Those observations raise two questions: (1) How does GPT-3.5 feedback differ from human feedback? (2) How does human/GPT-3.5 feedback impact response refinement? We address these questions in subsequent sections.
7.2 Compare Human and GPT-3.5 Feedback
To understand why adding human feedback is more beneficial than GPT-3.5 feedback, we analyze their differences through the following perspectives. (1) Refinement Success Rate: Percentage of satisfactory feedback-driven refinements. (2) Verbosity: Average word count of feedback. (3) Diversity: Percentage of unique words. (4) Grammar: Percentage of grammatical feedback sentences.888 We use Gramformer for grammar error checking: https://github.com/PrithivirajDamodaran/Gramformer.
Figure 2: Compare human and GPT-3.5 feedback.
Refinement GRD Fact. Help. Rel. Conf. Sat.
No feedback 39.16 90.35 83.48 98.10 100.00 76.50
Human FB 40.11 87.50 81.10 97.80 99.84 74.60
GPT-3.5 FB 32.77 81.50 90.20 98.40 99.84 79.50
Table 10: Quality of refinements with no/human/GPT-3.5 feedback. See Table 6 for abbreviation meanings.
In Figure 2, we show characteristics of human and GPT-3.5 feedback. Though GPT-3.5 feedback is lengthier and grammatically sound, it lacks the language diversity of human feedback. Upon manual examination, GPT-3.5 feedback is often general, whereas human feedback is direct and specific. See the Appendix for feedback examples.
7.3 Feedback Impact on Refinements
While GPT-3.5 feedback leads to more satisfactory refinements (see Figure 2), the performance of the resulting dialog system trained with these refinements falls short w.r.t. all our designed metrics compared to the system trained using human feedback-driven refinements as shown in Table 9. This disparity merits further scrutiny. Therefore, we also evaluate the refinement quality via designed metrics from §6.1.1, with results in Table 10. Refinements obtained using human feedback mainly stand out in groundedness and factuality. This aligns with the feedback clusters in Table 5 where over 40% of the feedback suggests the bot focus more on the search results; that is, focusing more on the search results will make the refinements more grounded, leading to a more grounded final system (see Table 9). Since language models are known to hallucinate regardless of their size (Ji et al., 2023; Li et al., 2023), grounding their generations to the documents is important to ensure factuality. Hence, groundedness of refinements plays an essential role in the performance of trained models.
7.4 Advantages of Human Feedback
We find that human feedback pinpoints issues more effectively than GPT-3.5 feedback. For example, when a response does not answer a question, GPT-3.5 will say that the response is unhelpful because it does not contain the information the user wants. In contrast, human feedback often provides specific hints from the search results, guiding the model towards a better response. Thus, despite GPT-3.5 producing seemingly informative feedback, it currently can’t match the nuance of human annotators.
8 Conclusion
In this paper, we present a framework that harnesses system-level NL feedback. By using a large set of instance-level feedback, we derive system-level feedback for refinement prompt engineering and metric design. We show the effectiveness of system-level feedback through two case studies: generating queries and formulating dialogue responses. We further combine system-level and instance-level feedback in the refinement data construction process, and observe that the resulting trained response generator makes considerable improvements versus using system-level feedback only. Finally, we explore the possibility of substituting instance-level human feedback with GPT-3.5 feedback. We find that human feedback stands out in capturing main issues, while GPT-3.5 feedback is lengthy and less focused.
9 Limitations
Due to the lack of publicly available natural language feedback datasets, our experiments were limited to the small-scale dialog system BB2, which does not represent the current state-of-the-art. We recognize that integrating more advanced models such as ChatGPT could yield further insights, presenting a promising direction for future research. As relevant datasets become more accessible, we look forward to exploring these possibilities.
Appendix A Appendix
A.1 Manual Efforts Required to Derive System-level Criteria
In our approach, the feedback grouping was a semi-automated process. Initially, we employed k-means clustering, utilizing the SimCSE encoder to categorize the feedback sentences. This clustering process was conducted once without human intervention. Then we employed a streamlined, non-iterative manual approach for cluster curation. Specifically, two domain experts independently reviewed 50 samples (each requires around 30s to read) from each of the 15 clusters (including clusters for queries and responses), requiring approximately 375 minutes per person for this phase. This was followed by a collaborative discussion to merge insights and remove duplicate clusters, amounting to an additional 60 minutes per expert. Thus, the total human effort amounted to approximately 14.5 person-hours.
A.2 Refinement with GPT-3.5
We instruct GPT-3.5 to generate query refinements and response refinements using carefully crafted prompts, as shown in Table 11 and Table 12. The corresponding ablation studies with full criteria text for query refinement and response refinement are shown in Table 13 and Table 14. The prompt for using both system-level feedback and instance-level feedback for response refinement is shown in Table 15.
Prompt for query refinement with GPT-3.5
[Criteria]
Below is the dialog context.
[Dialog Context]
Below is the bot’s unsatisfactory query.
[Original Query]
You should modify the original search query into the following:
Prompt for response refinement with GPT-3.5
[Criteria]
Below is the dialog context.
[Dialog Context]
Below is the bot’s unsatisfactory response.
[Original Response]
Below are some useful search results that you could use.
[Search Documents]
You should modify the original response into the following:
@@ A.1 Manual Efforts Required to Derive System-level Criteria
A.3 Evaluation with GPT-3.5
Since previous studies have demonstrated GPT-3’s capability in the evaluation of aspects such as factuality (Luo et al., 2023), helpfulness (Fu et al., 2023), relevance (Fu et al., 2023), etc. We use GPT-3.5 to evaluate the following perspectives with the help of ChatEval (Yuan et al., 2023).
Query Specificity